# Stylos: Multi-View 3D Stylization with Single-Forward Gaussian Splatting

### v1

@@ Abstract
We present Stylos, a single-forward 3D Gaussian framework for 3D style transfer that operates on unposed content, from a single image to a multi-view collection, conditioned on a separate reference style image. Stylos synthesizes a stylized 3D Gaussian scene without per-scene optimization or precomputed poses, achieving geometry-aware, view-consistent stylization that generalizes to unseen categories, scenes, and styles. At its core, Stylos adopts a Transformer backbone with two pathways: geometry predictions retain self-attention to preserve geometric fidelity, while style is injected via global cross-attention to enforce visual consistency across views. With the addition of a voxel-based 3D style loss that aligns aggregated scene features to style statistics, Stylos enforces view-consistent stylization while preserving geometry. Experiments across multiple datasets demonstrate that Stylos delivers high-quality zero-shot stylization, highlighting the effectiveness of global style–content coupling, the proposed 3D style loss, and the scalability of our framework from single view to large-scale multi-view settings. Our codes will be available at https://github.com/HanzhouLiu/StylOS.
@@ 1 Introduction
In contrast, we introduce Stylos (meaning pens in French), a single-forward framework for 3D style transfer that eliminates the need for per-scene optimization and precomputed camera parameters, and generalizes effectively to novel categories, scenes, and styles. Stylos employs a shared Transformer backbone with two pathways: content and style images are projected into a shared feature space, where content retains self-attention for geometric reasoning, and style is injected via global cross-attention to condition aesthetics. Notably, the global cross-attention also benefits the preservation of geometric details. Geometry-related attributes, such as depth, camera intrinsics, and extrinsics, are derived from backbone features, whereas style conditioning guides the prediction of color coefficients. These outputs are estimated through prediction heads that serve as the interface between feature space and the final Gaussian representation. Prior works in 2D style transfer typically rely on distribution alignment (Li et al., 2017), which operate on image statistics but lack explicit multi-view or 3D consistency. To address this, we explore alternative objectives and propose a voxel-based 3D style loss that aligns aggregated scene features with style statistics, providing stronger view-consistent stylization while preserving geometric fidelity.
@@ 2.1 Single-Forward 3D Reconstruction
Explicit-representation approaches to 3D stylization (Kato et al., 2018; Cao et al., 2020; Guo et al., 2021; Huang et al., 2021; Höllein et al., 2022; Mu et al., 2022) have shown promising results, but their reliance on accurate geometry limits applicability in complex real-world scenes. Implicit neural representations like NeRF (Mildenhall et al., 2020) improve geometric consistency, inspiring a range of stylization methods based on stylized view supervision or auxiliary networks (Nguyen-Phuoc et al., 2022; Zhang et al., 2022; Chiang et al., 2022; Huang et al., 2022). StyleRF (Liu et al., 2023) further enhances fidelity and enables zero-shot transfer by applying style transformations in radiance-field feature space.
More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as an efficient alternative to NeRF. StyleGaussian (Liu et al., 2024) extends 3DGS with efficient feature rendering to enable real-time stylization, while subsequent works refine 3DGS for specialized settings and efficiency improvements (Mei et al., 2024; Zhang et al., 2024; Yu et al., 2024; Galerne et al., 2025). In parallel, Styl3R (Wang et al., 2025b) disentangles geometry and style using separate encoders, but is restricted to 2–8 input views. In contrast, Stylos follows the shared-backbone paradigm of AdaIN (Huang & Belongie, 2017), leveraging VGGT for unified feature extraction, and scales from a single to hundreds of input views.
Underlying these architectures, style transfer objectives play a crucial role. Classical style transfer relies on Gram-matrix correlations (Gatys et al., 2016), while channel-wise statistics such as mean and variance (Li et al., 2017) offer efficient alternatives, forming the basis of AdaIN (Huang & Belongie, 2017). Extensions to video and multi-view stylization (Ruder et al., 2016; Gupta et al., 2017; Nguyen-Phuoc et al., 2022; Zhang et al., 2022) address temporal and cross-view consistency, and CLIP-based losses (Radford et al., 2021) introduce semantic alignment with text or image prompts. However, most objectives remain inherently 2D and cannot guarantee consistency across complex 3D scenes. To this end, we extend statistics-based style losses into voxel space, introducing a 3D-aware objective that aligns feature distributions after multi-view fusion. This formulation complements perceptual and CLIP-based losses, enforcing geometry-aware and view-consistent stylization.
@@ 3.1 Problem Formulation
where the scene representation $G=\{(p_{m},\,c_{m})\}_{m=1}^{M}$ is parameterized by $M$ anisotropic 3D Gaussians. Each Gaussian has geometry attributes $p_{m}=(\mu_{m},\,\alpha,\,r_{m},\,s_{m})$, comprising 3D position $\mu_{m}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}_{+}$, orientation quaternion $r_{m}\in\mathbb{R}^{4}$, and anisotropic scale $s_{m}\in\mathbb{R}^{3}$, as well as a style-dependent color embedding $c_{m}\in\mathbb{R}^{3\times(k+1)^{2}}$ represented with spherical harmonics of degree $k$. In addition, Stylos jointly predicts camera parameters $\{g_{i}\in\mathbb{R}^{9}\}_{i=1}^{N}$ for each input view, leveraging pose cues estimated by the VGGT backbone.
@@ 3.2 Network Architecture
Architecture Overview. At its core, Stylos employs a geometry backbone that alternates frame-wise and global attention to process multiple input views and infer geometry-related parameters such as positions, scales, orientations, and opacities. To enable stylization, we introduce a dedicated conditioning branch, the Style Aggregator, which fuses content and style features through cross-attention and predicts style-aware color embeddings for each Gaussian. In this way, geometry remains derived solely from the backbone, while style representation is conditioned on the style reference. We next detail the geometric backbone in Sec. 3.2.1, the style aggregation module in Sec. 3.2.2, and the prediction heads in Sec. 3.2.3.
@@ 3.2.2 Style Aggregator
The Style Aggregator extends the geometric backbone by adapting its attention layers from self-attention to cross-attention. Both content views and the style image are embedded with a shared patch encoder, ensuring that their tokens lie in a common feature space. In this design, content tokens serve as queries while style tokens provide keys and values, so that style information conditions the token representations later used for predicting Gaussian colors. Geometry-related parameters, in contrast, are still predicted from the backbone pathway and remain unaffected by style conditioning.
Formally, let $T\in\mathbb{R}^{B\times N\times K\times C}$ denote the content tokens from $N$ views, each with $K$ patches and $C$ channels, and let $S\in\mathbb{R}^{B\times K\times C}$ represent the style tokens. Let $T_{i}$ be the token set of the $i$-th view, and let $\mathrm{Attn}(Q,KV)$ denote a standard cross-attention operation. Depending on how queries are formed, we distinguish three style–content coupling mechanisms,
(Frame-only): $T^{\text{frame}}_{i}=\mathrm{Attn}(T_{i},S),$ $Q=T_{i},\;\;K=V=S,$ (1)
(Global-only): $T^{\text{global}}=\mathrm{Attn}(\{T_{i}\}_{i=1}^{N},S),$ $Q=\{T_{i}\}_{i=1}^{N},\;\;K=V=S,$ (2)
(Hybrid): $T^{\text{hybrid}}=[\,\{T^{\text{frame}}_{i}\}_{i=1}^{N}\,\|\,T^{\text{global}}\,],$ concatenation along channel dim. (3)
Frame-only injects style independently per view, global-only enforces a shared conditioning across all views, and hybrid concatenates the frame-level and global-level outputs along the channel dimension, yielding $2\times C$-dimensional features. In our design, Stylos adopts the global-only configuration, implemented with CrossBlock layers where $\{T_{i}\}_{i=1}^{N}$ serve as queries and $S$ provides keys and values. The resulting features are passed to the prediction heads described in Sec. 3.2.3.
@@ 3.2.3 Prediction Heads
Auxiliary Heads, Adapter, and Voxelization. We employ the existing VGGT camera head to estimate camera intrinsics and extrinsics, and a depth head to predict scene geometry cues, which are unprojected into 3D anchors for Gaussian placement (Kerbl et al., 2023; Ren et al., 2024). A Gaussian adapter then consolidates geometry and style outputs into a unified set of primitives $\{(p_{m},c_{m})\}_{m=1}^{M}$ for differentiable rendering. Finally, to reduce redundancy and balance density, we follow the voxelization step introduced by AnySplat (Jiang et al., 2025), nearby points are clustered within a discretized 3D grid and fused using confidence-aware weighting. This operation depends only on the unprojected 3D points and features, and is independent of camera parameters.
@@ 3.4 Style Losses
We denote by $F^{l}_{b,s}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the VGG feature map of the rendered (stylized) image for the $b$-th scene and $s$-th view at layer $l$, and by $S^{l}_{b}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the feature map of the style reference image. Building on the BN-statistics formulation (Li et al., 2017), which interprets channel-wise mean and variance as style descriptors, we progressively extend statistics matching beyond the standard 2D setting. This progression starts from independent image-level matching, moves to multi-view feature aggregation for enforcing cross-view consistency, and culminates in a voxel-space loss that directly constrains the fused 3D representation.
@@
Image-Level Style Loss. The simplest baseline aligns each rendered frame independently with the style reference,
$\mathcal{L}^{\mathrm{img}}_{\mathrm{sty}}=\frac{1}{BS}\sum_{b=1}^{B}\sum_{s=1}^{S}\sum_{l}\alpha_{l}\Bigl(\|\mu(F^{l}_{b,s})-\mu(S^{l}_{b})\|_{2}^{2}+\|\sigma(F^{l}_{b,s})-\sigma(S^{l}_{b})\|_{2}^{2}\Bigr).$ (4)
This encourages per-frame stylization but does not enforce multi-view consistency.
@@
Scene-Level Style Loss. To encourage consistent stylization across multiple views, we concatenate the per-view features $\{F^{l}_{b,s}\}_{s=1}^{S}$ along the spatial dimension to form $\widetilde{F}^{l}_{b}$, and compute statistics on this aggregated map,
@@
$\mathcal{L}^{\mathrm{scn}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l}\alpha_{l}\Bigl(\|\mu(\widetilde{F}^{l}_{b})-\mu(S^{l}_{b})\|_{2}^{2}+\|\sigma(\widetilde{F}^{l}_{b})-\sigma(S^{l}_{b})\|_{2}^{2}\Bigr).$ (5)
This step enforces cross-view consistency but still operates in 2D feature space.
@@
3D Style Loss. Finally, we fuse multi-view features into a voxel grid using differentiable unprojection, where features from different views are accumulated into spatial bins of a discretized 3D volume. Let $G^{l}_{b}$ denote the voxelized features for the $b$-th scene at layer $l$. We then compute style statistics directly in voxel space,
$\mathcal{L}^{\mathrm{3D}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l}\alpha_{l}\Bigl(\|\mu(G^{l}_{b})-\mu(S^{l}_{b})\|_{2}^{2}+\|\sigma(G^{l}_{b})-\sigma(S^{l}_{b})\|_{2}^{2}\Bigr).$ (6)
@@
By operating on voxelized features, this loss explicitly encodes geometry and enforces style consistency across both views and the underlying 3D structure.

### v2

@@ Abstract
We present Stylos, a single-forward 3D Gaussian framework for 3D style transfer that operates on unposed content, from a single image to a multi-view collection, conditioned on a separate reference style image. Stylos synthesizes a stylized 3D Gaussian scene without per-scene optimization or precomputed poses, achieving geometry-aware, view-consistent stylization that generalizes to unseen categories, scenes, and styles. At its core, Stylos adopts a Transformer backbone with two pathways: geometry predictions retain self-attention to preserve geometric fidelity, while style is injected via global cross-attention to enforce visual consistency across views. With the addition of a voxel-based 3D style loss that aligns aggregated scene features to style statistics, Stylos enforces view-consistent stylization while preserving geometry. Experiments across multiple datasets demonstrate that Stylos delivers high-quality zero-shot stylization, highlighting the effectiveness of global style-content coupling, the proposed 3D style loss, and the scalability of our framework from single view to large-scale multi-view settings. Our codes are available at https://github.com/HanzhouLiu/Stylos.
footnotetext: * Equal Contribution
@@ 1 Introduction
In contrast, we introduce Stylos (meaning pens in French), a single-forward framework for 3D style transfer that eliminates the need for per-scene optimization and precomputed camera parameters, and generalizes effectively to novel categories, scenes, and styles. Stylos employs a shared Transformer backbone with two pathways: content and style images are projected into a shared feature space, where content retains self-attention for geometric reasoning, and style is injected via global cross-attention to condition aesthetics. Notably, the global cross-attention also helps preserve fine visual details in the Gaussian splatting renderings. Geometry-related attributes, such as depth, camera intrinsics, and extrinsics, are derived from backbone features, whereas style conditioning guides the prediction of color coefficients. These outputs are estimated through prediction heads that serve as the interface between feature space and the final Gaussian representation. Prior works in 2D style transfer typically rely on distribution alignment (Li et al., 2017), which operate on image statistics but lack explicit multi-view or 3D consistency. To address this, we explore alternative objectives and propose a voxel-based 3D style loss that aligns aggregated scene features with style statistics, providing stronger view-consistent stylization while preserving geometric fidelity.
@@ 2.1 Single-Forward 3D Reconstruction
Explicit-representation approaches to 3D stylization (Kato et al., 2018; Cao et al., 2020; Guo et al., 2021; Huang et al., 2021; Höllein et al., 2022; Mu et al., 2022) have shown promising results, but their reliance on accurate geometry limits applicability in complex real-world scenes. Implicit neural representations like NeRF (Mildenhall et al., 2020) improve geometric consistency, inspiring a range of stylization methods based on stylized view supervision or auxiliary networks (Nguyen-Phuoc et al., 2022; Zhang et al., 2022; Chiang et al., 2022; Huang et al., 2022). StylizedNeRF (Huang et al., 2022) addresses the domain gap between style images and NeRF by introducing a mutual learning framework. StyleRF (Liu et al., 2023) further enables zero-shot transfer by applying style transformations in radiance-field feature space. However, training NeRF for scene reconstruction is typically computationally expensive, making NeRF-based 3D stylization far from real-time.
More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as an efficient alternative to NeRF for novel view synthesis. Subsequent works further refine 3DGS for efficient stylization (Mei et al., 2024; Zhang et al., 2025; Kovács et al., 2024; Yu et al., 2024; Galerne et al., 2025). StyleGaussian (Liu et al., 2024) extends 3DGS with efficient feature rendering to enable real-time zero-shot stylization. However, despite being significantly faster to optimize than NeRF, 3DGS-based approaches still require per-scene fitting, which poses a fundamental barrier to achieving truly real-time 3D stylization. Styl3R (Wang et al., 2025b), the closest contemporaneous related work to ours, introduces a feedforward framework for instant 3D stylized reconstruction. While effective and efficient, Styl3R is primarily designed for 2–8 input views and does not specifically target strong multi-view consistency among the rendered stylization results.
Underlying these architectures, style transfer objectives play a crucial role. Classical style transfer relies on Gram-matrix correlations (Gatys et al., 2016), while channel-wise statistics such as mean and variance (Li et al., 2017) offer efficient alternatives, forming the basis of AdaIN (Huang & Belongie, 2017). Extensions to video and multi-view stylization (Ruder et al., 2016; Gupta et al., 2017; Nguyen-Phuoc et al., 2022; Zhang et al., 2022) address temporal and cross-view consistency, and CLIP-based losses (Radford et al., 2021) introduce semantic alignment with text or image prompts. However, most objectives remain inherently 2D and cannot guarantee consistency across complex 3D scenes. To this end, we extend statistics-based style losses into voxel space, introducing a 3D-aware objective that aligns feature distributions after multi-view fusion.
@@ 3.1 Problem Formulation
where the scene representation $G=\{(p_{m},\,c_{m})\}_{m=1}^{M}$ is parameterized by $M$ anisotropic 3D Gaussians. Each Gaussian has geometry attributes $p_{m}=(\mu_{m},\,\alpha_{m},\,r_{m},\,s_{m})$, comprising 3D position $\mu_{m}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}^{+}$, orientation quaternion $r_{m}\in\mathbb{R}^{4}$, and anisotropic scale $s_{m}\in\mathbb{R}^{3}$, as well as a style-dependent color embedding $c_{m}\in\mathbb{R}^{3\times(k+1)^{2}}$ represented with spherical harmonics of degree $k$. In addition, Stylos jointly predicts camera parameters $\{g_{i}\in\mathbb{R}^{9}\}_{i=1}^{N}$ for each input view, leveraging pose cues estimated by the VGGT backbone.
@@ 3.2 Network Architecture
Stylos employs a geometry backbone that alternates frame-wise and global attention to process multiple input views and infer geometry-related parameters such as positions, scales, orientations, and opacities. To enable stylization, we introduce a dedicated conditioning branch, the Style Aggregator, which fuses content and style features through cross-attention and predicts style-aware color embeddings for each Gaussian. In this way, geometry remains derived solely from the backbone, while style representation is conditioned on the style reference. We next detail the geometric backbone in Sec. 3.2.1, the style aggregation module in Sec. 3.2.2, and the prediction heads in Sec. 3.2.3.
@@ 3.2.2 Style Aggregator
The Style Aggregator builds on the geometric backbone inherited from VGGT, which is composed of standard Transformer $\mathrm{Block}$ containing self-attention followed by feed-forward layers. To enable style conditioning, we replace this standard block with an adapted $\mathrm{CrossBlock}$, which inserts a cross-attention operation between the self-attention and MLP stages (Deng et al., 2022). In $\mathrm{CrossBlock}$, content tokens serve as queries ($\mathcal{Q}$) while style tokens provide keys and values ($\mathcal{KV}$). This structure allows the content representations to be refined by their own spatial context (via the preserved self-attention) while being explicitly conditioned on the target style (via the inserted cross-attention) before the final feed-forward projection.
Content–Style Coupling. Let $\mathcal{Q}_{b,v}\in\mathbb{R}^{L_{q}\times C}$ denote the content tokens extracted from the $v$-th view of the $b$-th sample, and let $\mathcal{KV}_{b}\in\mathbb{R}^{L_{kv}\times C}$ denote the style tokens. We use $\mathrm{CrossBlock}(\mathcal{Q},\mathcal{KV})$ to denote our block operator that updates $\mathcal{Q}$ using $\mathcal{KV}$ as key–value pairs. Depending on how the query set is formed, we consider two primary topological strategies.
(1) Frame cross attention. Each view independently interacts with the style tokens,
$\widetilde{\mathcal{Q}}_{b,v}=\mathrm{CrossBlock}(\mathcal{Q}_{b,v},\,\mathcal{KV}_{b}).$ (1)
Since the internal self-attention of the block operates only on $\mathcal{Q}_{b,v}$, this strategy preserves view-specific geometric structure but prevents information propagation between different views.
(2) Global cross attention. We first concatenate all views to obtain a global sequence $\mathcal{Q}_{b}^{\mathrm{global}}\in\mathbb{R}^{VL_{q}\times C}$ by $\mathcal{Q}_{b}^{\mathrm{global}}=\mathrm{Concat}_{v=1}^{V}\mathcal{Q}_{b,v}$, and perform the block operation simultaneously,
$\widetilde{\mathcal{Q}}^{\mathrm{global}}_{b}=\mathrm{CrossBlock}(\mathcal{Q}^{\mathrm{global}}_{b},\,\mathcal{KV}_{b}).$ (2)
We then reshape the updated tokens back to per-view tensors, $\widetilde{\mathcal{Q}}_{b,v}$, by splitting $\widetilde{\mathcal{Q}}^{\mathrm{global}}_{b}$ along the sequence dimension. This approach enables long-range reasoning: the internal self-attention ensures multi-view geometric consistency, while the cross-attention broadcasts style information globally.
(3) Hybrid cross attention. We first apply frame attention to refine each view independently as in Eq. 1, and then perform global attention by concatenating the resulting $\widehat{\mathcal{Q}}_{b,v}$ across views and applying the same operator as in Eq. 2. This hierarchical coupling combines per-view refinement with multi-view aggregation.
@@ 3.2.3 Prediction Heads
Auxiliary Heads, Adapter, and Voxelization. We employ the existing VGGT camera head to estimate camera intrinsics and extrinsics, and a depth head to predict scene geometry cues, which are unprojected into 3D anchors for Gaussian placement (Kerbl et al., 2023; Ren et al., 2025). A Gaussian adapter then consolidates geometry and style outputs into a unified set of primitives $\{(p_{m},c_{m})\}_{m=1}^{M}$ for differentiable rendering. Finally, to reduce redundancy and balance density, we follow the voxelization step introduced by AnySplat (Jiang et al., 2025), nearby points are clustered within a discretized 3D grid and fused using confidence-aware weighting. This operation depends only on the unprojected 3D points and features, and is independent of camera parameters.
@@ 3.3 Training Strategy
In our experiments, we set $\lambda_{\text{distill}}=1.0$, and use $\{{\lambda_{\text{style}},\lambda_{\text{cnt}},\lambda_{\text{clip}},\lambda_{\text{tv}}}\}=\{{1.0,0.1,1.0,10.0}\}$
@@ 3.4 Style Losses
We denote by $\mathcal{R}^{l}_{b,v}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the VGG feature map of the rendered image for the $b$-th scene and $v$-th view at layer $l$, and by $\mathcal{S}^{l}_{b}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the feature map of the single style image. Building on the Batch Normalization statistics method (Li et al., 2017), which interprets BN mean and variance as style descriptors, we progressively extend it matching beyond standard 2D settings. This progression starts from independent image-level matching, moves to multi-view feature aggregation that promotes global style coherence across views, and culminates in a voxel-space loss that directly constrains the fused 3D representation. Next, we formulate the three losses studied in this paper.
@@
Image-Level Style Loss. The simplest baseline aligns each rendered frame independently with the style reference, which encourages per-frame stylization but does not enforce multi-view consistency,
$\mathcal{L}^{\mathrm{img}}_{\mathrm{sty}}=\frac{1}{BV}\sum_{b=1}^{B}\sum_{v=1}^{V}\sum_{l=1}^{5}\alpha_{l}\Bigl(\|\mu(\mathcal{R}^{l}_{b,v})-\mu(\mathcal{S}^{l}_{b})\|_{2}^{2}+\|\sigma(\mathcal{R}^{l}_{b,s})-\sigma(\mathcal{S}^{l}_{b})\|_{2}^{2}\Bigr),$ (3)
where $\alpha_{l}$ denotes the weight at each level $l$, $\mu$ is the mean and $\sigma$ is the standard deviation.
@@
Scene-Level Style Loss. To introduce consistency into multi-view 3D stylization, we concatenate the per-view features $\{\mathcal{R}^{l}_{b,v}\}_{v=1}^{V}$ along the spatial dimension to form $\widetilde{\mathcal{R}}^{l}_{b}$, and calculate statistics on this aggregated map. However, this method still operates in 2D feature space and is computed by,
@@
$\mathcal{L}^{\mathrm{scn}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l=1}^{5}\alpha_{l}\Bigl(\|\mu(\widetilde{\mathcal{R}}^{l}_{b})-\mu(\mathcal{S}^{l}_{b})\|_{2}^{2}+\|\sigma(\widetilde{\mathcal{R}}^{l}_{b})-\sigma(\mathcal{S}^{l}_{b})\|_{2}^{2}\Bigr).$ (4)
Algorithm 1 3D Voxel-based AdaIN Style Loss
0: Rendered features $\mathcal{R}$, Style features $\mathcal{S}$, 3D points $\mathcal{P}$, Confidence $\mathcal{C}$, Valid mask $\mathcal{M}$
0: 3D style loss $\mathcal{L}_{sty}^{\text{3D}}$
1: Initialize $\mathcal{L}_{sty}^{3D}\leftarrow 0$
2:for $b=1$ to $B$ do
3:  for $l=1$ to 55 do
4:   Resize $(\mathcal{C,M,P})$ to match the resolution of $\mathcal{R}^{l}_{b}$
5:   // Next, fuse all views into a single voxel grid
6:   $\mathcal{G}_{b}^{l}\leftarrow\text{VoxelizeAndFuse}\left(\{\mathcal{R}^{l}_{b,v}\}_{v=1}^{V},\{\mathcal{P}_{b,v}\}_{v=1}^{V},\{\mathcal{C}_{b,v}\}_{v=1}^{V},\{\mathcal{M}_{b,v}\}_{v=1}^{V}\right)$
7:   // Next, compute the BN statistics (mean and standard deviation) of two feature maps
8:   $(\mu_{g},\sigma_{g})\leftarrow\text{MeanStd}(\mathcal{G}_{b}^{l})$, $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(\mathcal{S}_{b}^{l})$
9:   $\mathcal{L}_{sty}^{3D}\leftarrow\mathcal{L}_{sty}^{3D}+\alpha_{l}\|\mu_{g}-\mu_{s}\|_{2}^{2}+\alpha_{l}\|\sigma_{g}-\sigma_{s}\|_{2}^{2}$
10:  end for
11:end for
12:return $\mathcal{L}_{sty}^{3D}=\lambda_{sty}\mathcal{L}_{sty}^{3D}$
@@
Voxel-level 3D Style Loss. Finally, to further enforce multi-view consistency, we fuse multi-view features into a voxel grid using differentiable unprojection, where features from different views are accumulated into spatial bins of a discretized 3D volume. Let $\mathcal{G}^{l}_{b}$ denote the voxelized features for the $b$-th scene at layer $l$. We then compute style statistics directly in voxel space,
$\mathcal{L}^{\mathrm{3D}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l=1}^{5}\alpha_{l}\Bigl(\|\mu(\mathcal{G}^{l}_{b})-\mu(\mathcal{S}^{l}_{b})\|_{2}^{2}+\|\sigma(\mathcal{G}^{l}_{b})-\sigma(\mathcal{S}^{l}_{b})\|_{2}^{2}\Bigr).$ (5)
@@
By operating on voxelized features, this loss explicitly encodes geometry and enforces style consistency across both views and the underlying 3D structure. We provide pseudo codes in Algorithm 1.
4 Experiment
Datasets. We evaluate cross-category generalization on the CO3D (Reizenstein et al., 2021) dataset by training on 17 categories and testing on 3 held-out ones, and cross-scene generalization by training on the full DL3DV-10K (Ling et al., 2024) and testing on Tanks & Temples (Knapitsch et al., 2017). Style images are provided by WikiArt (WikiArt, 2025) and DELAUNAY (Gontier et al., 2023), with 50 diverse style images reserved as unseen styles that are never used during training.
Baselines. We compare Stylos with recent 3D stylization models. The baselines include, (1) a per-scene training and zero-shot method StyleGaussian (Liu et al., 2024), (2) per-scene and per-style optimization approaches G-Style (Kovács et al., 2024), StylizedGS (Zhang et al., 2025), and SGSST (Galerne et al., 2025), and (3) the closest related work Styl3R (Wang et al., 2025b).
Evaluation Metrics. Our evaluation covers three aspects. (1) To assess geometry reconstruction in the style-free setting, we report PSNR, SSIM, and LPIPS (Zhang et al., 2018) between content views and predictions, conditioned on the original first frame of each scene as the style input. (2) To measure stylization consistency, following prior work (Chiang et al., 2022), we compute LPIPS and RMSE in both short-range and long-range settings. (3) To further evaluate stylization quality, we report ArtScore (Chen et al., 2024), a recent metric specifically designed for reference-free evaluation of artness in generated images, and ArtFID (Wright & Ommer, 2022; Chung et al., 2024).
4.1 Ablation Study